Dexter has 36 pets. He has 3 times as many fish as mice and half as many cats as mice. How many of each pet does he have?
Answer by checkley77(12844)

F=3M
C=M/2
F+C+M=36
3M+M/2+M=36
4M+.5M=36
4.5M=36
M=36/4.5
M=8 MICE.
F=3*8=24 FISH.
C=8/2=4 CATS.
PROOF.
24+4+8=36
36=36